How to get 100 by using digits from 1 to 9. Rules-
Only addition can be used.......
every digit should and can used only once and compulsory to use every digit.........
pairs can be made like 2 and 3 digits can be used as 23 or 32........
Fractions can be used like 2 and 5 digits can be written as 2/5 or 5/2...........

Answer by Edwin McCravy(20054)

There are loads of solutions:

12 + 75 + 6/3 + 8/4 + 9 = 100
12 + 75 + 8/4 + 6/3 + 9 = 100
12 + 79 + 6/3 + 8/4 + 5 = 100
12 + 79 + 8/4 + 6/3 + 5 = 100
14 + 75 + 2/3 + 8/6 + 9 = 100
14 + 75 + 3/2 + 9/6 + 8 = 100
14 + 75 + 9/6 + 3/2 + 8 = 100
14 + 78 + 3/2 + 9/6 + 5 = 100
14 + 78 + 9/6 + 3/2 + 5 = 100
14 + 79 + 2/3 + 8/6 + 5 = 100
15 + 72 + 6/3 + 8/4 + 9 = 100
15 + 72 + 8/4 + 6/3 + 9 = 100
15 + 74 + 2/3 + 8/6 + 9 = 100
15 + 74 + 3/2 + 9/6 + 8 = 100
15 + 74 + 9/6 + 3/2 + 8 = 100
15 + 78 + 3/2 + 9/6 + 4 = 100
15 + 78 + 9/6 + 3/2 + 4 = 100
15 + 79 + 2/3 + 8/6 + 4 = 100
15 + 79 + 6/3 + 8/4 + 2 = 100
15 + 79 + 8/4 + 6/3 + 2 = 100
18 + 74 + 3/2 + 9/6 + 5 = 100
18 + 74 + 9/6 + 3/2 + 5 = 100
18 + 75 + 3/2 + 9/6 + 4 = 100
18 + 75 + 9/6 + 3/2 + 4 = 100
19 + 72 + 6/3 + 8/4 + 5 = 100
19 + 72 + 8/4 + 6/3 + 5 = 100
19 + 74 + 2/3 + 8/6 + 5 = 100
19 + 75 + 2/3 + 8/6 + 4 = 100
19 + 75 + 6/3 + 8/4 + 2 = 100
19 + 75 + 8/4 + 6/3 + 2 = 100
24 + 57 + 6/3 + 8/1 + 9 = 100
24 + 57 + 6/3 + 9/1 + 8 = 100
24 + 58 + 6/3 + 7/1 + 9 = 100
24 + 58 + 6/3 + 9/1 + 7 = 100
24 + 59 + 6/3 + 7/1 + 8 = 100
24 + 59 + 6/3 + 8/1 + 7 = 100
27 + 54 + 6/3 + 8/1 + 9 = 100
27 + 54 + 6/3 + 9/1 + 8 = 100
27 + 58 + 6/3 + 4/1 + 9 = 100
27 + 58 + 6/3 + 9/1 + 4 = 100
27 + 59 + 6/3 + 4/1 + 8 = 100
27 + 59 + 6/3 + 8/1 + 4 = 100
28 + 54 + 6/3 + 7/1 + 9 = 100
28 + 54 + 6/3 + 9/1 + 7 = 100
28 + 57 + 6/3 + 4/1 + 9 = 100
28 + 57 + 6/3 + 9/1 + 4 = 100
28 + 59 + 6/3 + 4/1 + 7 = 100
28 + 59 + 6/3 + 7/1 + 4 = 100
29 + 54 + 6/3 + 7/1 + 8 = 100
29 + 54 + 6/3 + 8/1 + 7 = 100
29 + 57 + 6/3 + 4/1 + 8 = 100
29 + 57 + 6/3 + 8/1 + 4 = 100
29 + 58 + 6/3 + 4/1 + 7 = 100
29 + 58 + 6/3 + 7/1 + 4 = 100
31 + 56 + 4/8 + 7/2 + 9 = 100
31 + 59 + 4/8 + 7/2 + 6 = 100
36 + 51 + 4/8 + 7/2 + 9 = 100
36 + 59 + 4/8 + 7/2 + 1 = 100
39 + 51 + 4/8 + 7/2 + 6 = 100
39 + 56 + 4/8 + 7/2 + 1 = 100

Edwin